If f(x) = 8 – 10x and g(x) = 5x + 4, what is the value of (fg)(–2)?
Answer by Leaf W.(135)

(fg)(-2) essentially means f(g(-2)) -- in other words, you need to find g(-2), and then plug that value into the function f.
*
g(-2) = 5(-2) + 4 = -10 + 4 = -6
*
f(g(-2)) = f(-6) = 8 - 10(-6) = 8 + 60 = 68
*
Thus, the value of (fg)(-2) is 68.